OpenStudy (anonymous):

What are the solutions for x in the proportion 2 times x plus 3 all over x equals 2 times x all over 2.?

OpenStudy (anonymous):

somebody help

OpenStudy (diyadiya):

\[\frac{x}{2x} + \frac{3}{x} = \frac{2x}{2}\] is this your question?

OpenStudy (diyadiya):

@HubbaHubba Is that your question?

OpenStudy (diyadiya):

\[\frac{2x+3}{x} = \frac{2x}{2}\] First do cross multiplication \[2 (2x+3) = x(2x)\] is this clear?